OpenStudy (cookie_2046):

Fiona has a jar filled with 1050 coins made of quarters and nickels the total value of the coins in 23.10 how many quarters and how many nickels are in the jar? a. 88 quarters; 62 nickels b. 80 quarters; 70 nickels c. 75 quarters; 75 nickels d. 62 quarters; 75 nickels

OpenStudy (anonymous):

q+n=1050 .25q+.05n=23.10 solve that system of equations to get the values of q and n.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

\[23.10 = x*.25 +y*.05\] \[x + y =1050\] solve the system of equations using elimination or substitution

OpenStudy (misty1212):

i would say we need a variable for the quarters, say \(Q\) is the number of quarters then since the total amount of coins is \(1050\) the number of nickles has to be \(1050-Q\) that means \[25!+5\times (1050-Q)=2310\] easier to work with pennies than dollars, no decimals

OpenStudy (misty1212):

oops i made a typo, meant \[25 Q+5\times (1050-Q)=2310\]

OpenStudy (misty1212):

you got this dear or no?